OpenStudy (anonymous):

The average of five integers is 20. The ave of the three middle integers is 8. Wat is the ave. of the first and last integers?

OpenStudy (unklerhaukus):

the average of x : \(\langle x\rangle=\frac1N\sum\limits_{i=1}^Nx_i\)

OpenStudy (anonymous):

what??? O_O

OpenStudy (unklerhaukus):

\[=\frac{x_1+x_2+\cdots+x_N}{N}\]

OpenStudy (anonymous):

\[(a+b+c+d+e)=100\] \[(b+c+d)=24\] \[a+e=100-24=76\] \[(a+e)/2=76/2=38\]

OpenStudy (anonymous):

i tried 7,8,9,18,18. Not correct. and also 6,6,12,18,18.

OpenStudy (anonymous):

Thanks! :D